OpenStudy (anonymous):

find the center and radius of the circle: x^2+y^2+10x-14y=-10

OpenStudy (anonymous):

we haven't talked about these since the first week of the semester, and yet we are expected to remember it on the final tomorrow...

hartnn (hartnn):

do u know how to complete the square ? and general equation of circle ?

OpenStudy (anonymous):

have to group the x and y terms together and then complete the square

OpenStudy (anonymous):

ah, I do x^2+10x+___ and then the same for the y's and complete both squares?

hartnn (hartnn):

exactly!

hartnn (hartnn):

and general equation of circle ?

OpenStudy (anonymous):

ok so now I have (x^2+10x+25)+(y^2-14y+49)

OpenStudy (anonymous):

whats next?

OpenStudy (anonymous):

oh and =-10

hartnn (hartnn):

the general equation of circle is \((x-h)^2+(y-k)^2=r^2\)

hartnn (hartnn):

where h,k is centre and r is radius

hartnn (hartnn):

bring it in that form

hartnn (hartnn):

and the constant is not just -10

OpenStudy (anonymous):

ah, so I pull the 25 and 49 to the other side?

hartnn (hartnn):

you added 25 right ? to balance the equation, you need to subtract 25 also

hartnn (hartnn):

x^2+y^2+10x-14y + 25 +49 -25 -49=-10

hartnn (hartnn):

and bring all constants to right side.

OpenStudy (anonymous):

ok, I'm following along on paper

hartnn (hartnn):

could you bring it in this form ? \((x-h)^2+(y-k)^2=r^2\)

OpenStudy (anonymous):

(x+5)^2 +(y-7)^2=64

hartnn (hartnn):

absolutely correct so whats h,k = ?

OpenStudy (anonymous):

-5,7

OpenStudy (anonymous):

radius is 8?

hartnn (hartnn):

both correct :)

OpenStudy (anonymous):

awesome, I vaguely remember how to do it haha. thanks for the refresher

hartnn (hartnn):

welcome ^_^